Calculate the de broglie wavelength associated with an electron moving with velocity of 1/10 th the velocity of the light
1
Expert's answer
2020-10-28T08:41:49-0400

Velocity of the electron "v = \frac{1}{10} \times c"

"v = \frac{1}{10} \times 3 \times 10^8 = 3 \times 10^7 \; m/s"

Mass of electron "m = 9.1 \times 10^{-31} \;kg"

De Broglie equation:

"λ = \frac{h}{p} = \frac{h}{mv}"

h – Plank’s constant

"h = 6.62 \times 10^{-34} \;Js"

p – momentum

v – velocity of electron

"λ = \frac{6.62 \times 10^{-34}}{9.1 \times 10^{-31} \times 3 \times 10^{7}}"

"λ = 2.24 \times 10^{-9}"

"10^{-9} \;m = nm"

"λ = 2.24 \;nm"